OpenStudy (anonymous):

Please help!! Bart and April are purchasing a house with a 15-year, 2/1 ARM for $315,000 at 5.85% with a 2/8 cap structure. What will the difference in payments be from year 2 to year 3?

OpenStudy (amistre64):

i worked at developing a formula that would incorporate Payment in terms of Balances \[B_c=B_ok^{ct}-B_ok^{nt}(\frac{1-k^{ct}}{1-k^{nt}})\] \[B_c=B_ok^{ct}(1-k^{(n-c)t}~(\frac{1-k^{ct}}{1-k^{nt}}))\] and i believe i came up with this monstrocity to test out :)

OpenStudy (amistre64):

my ti83 will be put to good use now

OpenStudy (amistre64):

to many buttons to press :) lets try the wolf

OpenStudy (amistre64):

the wolf gives me a balance after 2 years of: 287140.2061 which we run thru a payment calculation

OpenStudy (amistre64):

287140.2061(1+.0785/12)^(13*12)(1-(1+.0785/12))/(1-(1+.0785/12)^(13*12)) the wolf say 2942 with that one .... is that an option by chance?

OpenStudy (amistre64):

difference in payments ....

OpenStudy (amistre64):

the begining payments i get as 2632 2942 -2632 ------ 310 about

OpenStudy (anonymous):

thanks so much!

OpenStudy (amistre64):

youre welcome, and excel helped out alot ;)